bannytoslimey:

The number 1,078 is to be written as a product of its prime factors. Which one of the following is correct? A 7 × 11 × 14 B 2 × 72 × 11 C 72 × 22 D 23 × 33 × 5

aidanlord:

do yo know how to do thid

aidanlord:

this*

personification:

breh tis A

jhonyy9:

how you factorize this number of 1078 ?

jhonyy9:

like a first step you see 1078 is an even number so divide it by 2 and will get 1078/2 = 539 539 can divide it by 7 and get 539/7 = 77 77 divide by 7 and get 77/7 = 11 and now from these above wrote try get the right answer

personification:

hey uh, the only set of numbers that equals 1078 is a

personification:

do da math